# Prediction of Energy Resolution in the JUNO Experiment

### v1

@@ 4.5 Determination of the LS absolute light yield
To ensure consistency between the JUNO and Daya Bay simulations, several modifications are made to the Daya Bay detector simulation, including:
@@
  • The same quenching effect model and the $kB$ parameter discussed in Sec. 4.2.2 are employed.
  • The LS RI and the improved Cherenkov process are the same.
  • The LS optical properties, such as the emission spectrum, quantum yield, scattering length, and time profiles, are assumed to be identical. The shape of the LS absorption spectrum is the same, however the absolute absorption lengths differ. In Daya Bay, the absorption length is 27 m at 430 nm, as measured in the experiment.
  • In Daya Bay, each PMT PDE is considered to be the measured quantum efficiency (QE), assuming a 100% collection efficiency (CE). In JUNO, both QE and CE are taken into account, and their product, PDE at normal incidence, is constrained by the PMT mass testing data [17]. The same PMT optical model (more details in Section 4.6) is used to describe the PMT reflection and angular responses. However, different optical properties of the photocathode are applied for the 8-inch PMTs in Daya Bay and the 20-inch PMTs in JUNO.
@@
    Fix the scintillation light yield $Y$ of the LS and tune $f_{C}$ in the Daya Bay simulation to match the LS energy non-linearity curve [46]. Figure 8a shows that good agreement can be achieved between the simulation and calibration data in Daya Bay.
@@
(b) 60Co energy spectrum at the detector center from the Daya Bay simulation before (blue) and after (red) $Y$ and $f_{C}$ scaling. The visible energy spectrum is fitted with a Crystal ball function, with the dashed line representing the mean value of main peak of the simulated 60Co energy spectrum.
Figure 8: Comparisons of the LS energy non-linearity and the energy scale curves between the Daya Bay simulation and experimental data with the newly constrained parameters of $Y$ and $f_{C}$.
$Y$ [/MeV] $kB\,[\textrm{g}/\textrm{cm}^{2}/\textrm{MeV}$] $f_{C}$ Production cut [mm]
@@
Table 1: A summary of a few parameters used in this work to predict the energy resolution in JUNO.
4.6 PMT’s photon detection efficiency and optical model
@@ 4.6.1 LPMTs
The PDE responses of LPMTs in JUNO are determined using the results from the LPMT mass testing setups [17] and the developed PMT optical model [18]. Two mass testing setups, the scanning station and the container system, are used for LPMT acceptance tests and performance evaluation. The container system evaluates the characteristics of each LPMT, including PDE, DCR, gain, and features of SPE A large area pulsed light source with a central wavelength of 420 nm is used to illuminate the entire photocathode of the LPMTs. The scanning station performs more detailed characterizations for LPMTs using 7 LED sources deployed along the longitude. The light beam from each LED is approximately perpendicular to the PMT surface and has a diameter of about 5 mm with a central wavelength of 420 nm. By rotating the PMTs, the whole photocathode can be scanned. However, even if only about 5% of JUNO LPMTs are tested following this full scan procedure, the results are representative of the total JUNO LPMT production.
The number of PMT’s detected photons is assumed to follow a Poisson distribution. The average value measured with the container system is converted to the PDE defined by the scanning station system for all measured PMTs, because the light intensities in this system are calibrated by a reference PMT. The conversion coefficients are determined by comparing the PDEs measured by both the scanning station and the container system for the three different types of LPMTs, including HPK dynode-PMT, NNVT MCP-PMT, and NNVT HQE MCP-PMT. The PDE given by the scanning station is defined as the averaged value across the PMT area, and it is determined from PDEs measured by the 7 LEDs and their surface area weights. The area weights, calculated based on the respective positions on the photocathode of the 7 LEDs, are dependent on the LPMT types (HPK and NNVT), and are summarized in Table 2.
@@
In detector simulation, the position-dependence of the PDE along the latitude of the photocathode is modeled using the CE curves. The PDE is considered as a product of the QE and the CE. Figure 9a shows the CE as a function of the polar angle in the PMT’s local coordinates for the NNVT MCP-PMT (blue), NNVT HQE MCP-PMT (cyan), and HPK dynode-PMT (red). The CE curves are obtained by averaging the PDEs measured by the 7 LEDs along the zenith angles in the scanning station. The maximum CE values are set to 100% for the NNVT MCP-PMT and 93% for the HPK dynode-PMT, indicated by the electrostatic simulation results from HPK and NNVT private communications.
@@
where $i$ represents the index of the LED in the scanning station and $w_{i}$ denotes the surface area weight of the $i$-th LED as summarized in Table 2. This calculation is performed for each LPMT in the CD to ensure that the product of QE and CE is consistent with the measured PDE, as published in [17]. The QE spectral responses are also implemented in the simulation based on laboratory measurements, as shown in Fig. 9b. The NNVT MCP-PMT (red) and NNVT HQE MCP-PMT (blue) have identical QE spectral responses, while the HPK PMTs (violet) have a different response at higher wavelength ($\lambda>500\,\mathrm{nm}$). It is assumed that LPMTs of the same type share the same CE curve and QE spectral response in the simulation.
Figure 9: (a) Angular dependence of collection efficiency for NNVT MCP-PMT (red), NNVT HQE MCP-PMT (blue), and HPK dynode-PMT (violet); (b) Wavelength dependence of quantum efficiency for NNVT MCP-PMT (red), NNVT HQE MCP-PMT (blue), and HPK dynode-PMT (violet).
The PMT optical model [18] takes into account the angle of incidence (AOI) dependence of the PDE, as well as the reflections on the photocathode and the optical processes inside the PMTs. In this model, the PMT window is treated as a multi-layer optical stack, consisting of water, PMT glass, anti-reflective coating, photocathode, and vacuum layers from outside to inside. The anti-reflective coating and photocathode are considered coherent layers due to their comparable thicknesses with the light wavelength. The model incorporates the light interference effect and the multiple reflections between adjacent boundaries using the transfer matrix method.
The RI $n$, extinction coefficient $k$, and thickness $d$ of the anti-reflective coating and photocathode are determined in the wavelength range of 390 nm to 500 nm by analyzing the reflectance data of NNVT MCP-PMT, NNVT HQE MCP-PMT, and HPK PMT immersed in the LAB liquid. The optical properties of the other components inside the PMTs, such as dynode (MCP), supporting structure, and aluminum film, are constrained by the QE data. These parameters are available in [18] and are directly used in this work. With these inputs, the PMT optical model calculates both the reflectance and absorbance for a given photon and AOI, assuming uniformity for the anti-reflective coating and the photocathode. The reflected photons are then propagated by GEANT4. The absorbance at a specific AOI is converted to the QE using the QE calculated above and absorbance information at normal incidence. This optical model has been integrated into the detector simulation and used in this work. It has also been applied in the Daya Bay simulation, using the optical parameters obtained from the reflectance and QE data of an 8-inch Daya Bay PMT.
@@ 5.1.1 LPMT contribution
When converting a SPE from the detector simulation into a pulse in the electronics simulation, the amplitude of the pulse is modeled by a combined distribution of a Gaussian and an exponential function. The SPE charge resolution of each PMT, obtained from PMT mass testing system, is directly used as the sigma of the Gaussian component in the simulation. The exponential component exhibits distinct behaviors between HPK LPMTs and NNVT LPMTs. For the former, the contribution from the exponential component is found to be 1% with an average amplitude of 1.1 PE, resulting in a SPE relative variance of 0.4. While for the latter, the exponential component describes the large signal observed in the measured charge spectrum of MCP-PMT (blue line in Fig. 10a) with an average amplitude of 2.2 PE, resulting in a SPE relative variance of 0.7. The ratio of large signals in the SPE spectrum of MCP-PMT, as a function of the zenith angle on the photocathode, has been characterized by laboratory measurements, as shown in Fig. 10b. Good agreement between Monte Carlo simulation and experimental data is achieved, illustrated in Fig. 10a.
@@
Figure 10: (a) The measured charge spectrum of NNVT MCP-PMT exhibits a long tail, which is described by an exponential distribution in the Monte Carlo simulation. Good agreement between data and Monte Carlo is achieved. (b) The ratio of the large signals depends on the positions on the photocathode, and is extracted from experimental data.
The transit time and TTS are also implemented for both HPK and NNVT LPMTs. Since these quantities are only measured for $\sim 25\%$ of LPMTs [17], a random sampling based on the distributions is used to obtain transit time and TTS values for other LPMTs in the CD. The transit time and TTS values are position-dependent on the photocathode, and this effect is also considered in the electronics simulation based on dedicated measurements made during the PMT mass testing.
The after-pulses of LPMTs are modeled based on the results published in [59]. Since the earliest after-pulse component occurs at about 1 $\mu$s after the primary signals, the overlap with the signals is expected to be small and thus has a negligible impact on the energy resolution. The non-linearity effect of LPMTs is also taken into account in the electronics simulation. However, it primarily affects high-energy events rather than IBD ones.
@@ 5.1.2 Electronics contribution
In the LPMTs electronics simulation, the SPE average waveform is obtained using the PMT testing data for both NNVT and HPK LPMTs. This waveform serves as a template in the electronics simulation to model the SPE response. The amplitude of the waveform is determined by the gain and SPE charge resolution discussed earlier.
To model the overshoot effect, the same model as that used in the Daya Bay experiment [60] is adopted, but with parameters tuned specifically for the JUNO experiment. The overshoot model is parameterized using an exponential function plus a Gaussian function. To capture the onset of the overshoot, the exponential function is multiplied by a Fermi function. The maximum amplitude of the overshoot is set to be 1% of the primary pulse.
White noise is applied to the raw waveform by sampling a Gaussian function with a mean of 10% of the SPE amplitude. This value is based on the requirements of the JUNO experiment. The digitization process is also modeled in the simulation, taking into account the effects of the front-end analog-to-digital converter (FADC) resolution, the high-gain and low-gain amplifications of the SPE signal, as well as the baseline offset. The simulation also considers the non-linearity effects of the electronics. However, it is anticipated that these effects will have a negligible impact on the energy resolution for positrons coming from IBD events.
Overall, the LPMT electronics simulation in the CD takes into account various aspects of the LPMT response, including the SPE waveform, overshoot, white noise, digitization, and non-linearity effects. These simulations aim to accurately model the electrical responses of the LPMTs and the readout electronics, and their impact on the energy resolution of the detector.
@@ 5.2 SPMT electronics simulation
The SPMT electronics simulation aims to accurately model the response of SPMTs and the readout electronics. For each SPMT, a DCR value is randomly sampled and assigned based on the DCR distribution obtained from the SPMT mass production measurements [40]. The mean values of the TTS (1.6 ns) and the SPE charge resolution (33%) are shared by all SPMTs.
The key features of the CatiROC ASIC are implemented in the simulation based on the measurements in Ref. [28], including the two types of dead times, the 70 ns trigger dead time and the 9.3 $\mu$s digitization dead time. In cases where a hit is not triggered due to dead time, a fraction of its charge can still be calculated. This fraction is called the charge acceptance and is parameterized based on the time difference with the previous hit. In addition, the charge in picocoulombs is converted into an ADC unit, with two conversion coefficients used responsible for high gain (less than 10 PE) and low gain (more than 10 PE).
@@ 6.1 Reconstruction and calibration of PMT waveforms
As described in Sec. 5, the PMT together with the electronics system will convert the photons to waveforms for each LPMT and to the pairs of charge and time for each SPMT. For each LPMT waveform, the charge and hit time of every pulse are reconstructed using the deconvolution algorithm [61], which exhibits smaller charge non-linearity compared to other algorithms. Meanwhile, the hit time of the first pulse, referred to as the first hit time, is of particular importance. A dedicated algorithm with a linear fit of the rising edge of the first pulse is developed to achieve a more accurate first hit time and reduce its charge dependence. Given that the charge and hit time information of PMTs are the inputs to the vertex and energy reconstruction in Sec. 6.2, the performance of the waveform reconstruction will also contribute to the energy resolution. Moreover, the CD has about O(104) PMTs and their characteristics may differ. Thus the reconstructed charge and hit time of each PMT must be calibrated to account for the different PMT parameters such as gain, TTS, relative PDE, and DCR. A comprehensive calibration strategy was developed in Ref. [24] to extract these parameters for all the PMTs and continuously monitor their time dependence.

### v2

@@ 4.5 Determination of the LS absolute light yield
To ensure consistency between the JUNO and Daya Bay simulations, several modifications were made to the Daya Bay detector simulation as follows:
@@
  • The same quenching effect model and $kB$ parameter discussed in Sec. 4.2.2 are employed.
  • The LS refractive index and improved Cherenkov process are the same.
  • The LS optical properties, such as the emission spectrum, quantum yield, scattering length, and time profiles, are assumed to be identical. The shape of the LS absorption spectrum is the same; however, the absolute absorption lengths differ. In Daya Bay, the absorption length is 27 m at 430 nm, as measured in the experiment.
  • In Daya Bay, each PMT PDE is considered to be the measured quantum efficiency (QE), assuming a 100% collection efficiency (CE). In JUNO, both QE and CE are taken into account, and their product, PDE at normal incidence, is constrained by the PMT mass testing data [21]. The same PMT optical model (more details in Section 4.6) is used to describe the PMT reflection and angular responses. However, different optical properties of the photocathode are applied for the 8-inch PMTs in Daya Bay and 20-inch PMTs in JUNO.
@@
    Fix the scintillation light yield $Y$ of the LS and tune $f_{C}$ in the Daya Bay simulation to match the gamma energy non-linearity curve in [52]. By tuning $f_{C}$, the ratio of scintillation photons to Cherenkov photons, which determines the shape of the energy non-linearity curve, can be adjusted. Figure 8a shows that good agreement can be achieved between the simulation and calibration data in Daya Bay.
@@
(b) 60Co energy spectrum at the detector center from the Daya Bay simulation before (blue shaded area) and after (red shaded area) $Y$ and $f_{C}$ scaling. The visible energy spectrum is fitted with a Crystal ball function combined with a Gaussian distribution (solid line), while the dashed line denotes the mean value of main peak of the simulated 60Co energy spectrum.
Figure 8: Comparisons of the LS energy non-linearity and energy scale curves between the Daya Bay simulation and experimental data with the newly constrained parameters of $Y$ and $f_{C}$.
$Y$ [/MeV] $kB~{}[\textrm{g}/\textrm{cm}^{2}/\textrm{MeV}$] $f_{C}$ Production cut [mm]
@@
Table 1: Summary of a few parameters used in this work to predict the energy resolution in JUNO.
4.6 PMT photon detection efficiency and optical model
@@ 4.6.1 LPMTs
The PDE responses of LPMTs in JUNO are determined using the results from the LPMT mass testing setups [21] and the developed PMT optical model [22]. Two mass testing setups, the scanning station and container system, are used for LPMT acceptance tests and performance evaluation. The container system evaluates the characteristics of each LPMT, including PDE, DCR, gain, and features of SPE. A large area pulsed light source with a central wavelength of 420 nm is used to illuminate the entire photocathode of the LPMTs. The scanning station performs more detailed characterizations for LPMTs using 7 LED sources deployed along the longitude. The light beam from each LED is approximately perpendicular to the PMT surface and has a diameter of approximately 5 mm with a central wavelength of 420 nm. By rotating the PMTs, the whole photocathode can be scanned. However, even if only approximately 5% of JUNO LPMTs are tested following this full scan procedure, the results are representative of the total JUNO LPMT production.
The number of PMT detected photons is assumed to follow a Poisson distribution. The average value measured with the container system is converted to the PDE defined by the scanning station system for all measured PMTs, because the light intensities in this system are calibrated by a reference PMT. The conversion coefficients are determined by comparing the PDEs measured by both the scanning station and container system for the three different types of LPMTs: HPK dynode-PMT, NNVT MCP-PMT, and NNVT HQE MCP-PMT. The PDE given by the scanning station is defined as the averaged value across the PMT area, and it is determined from PDEs measured by the 7 LEDs and their surface area weights. The area weights, calculated based on the respective positions on the photocathode of the 7 LEDs, are dependent on the LPMT types (HPK and NNVT) and are summarized in Table 2.
@@
In the detector simulation, the position dependence of the PDE along the latitude of the photocathode is modeled using the CE curves. The PDE is considered as a product of the QE and CE. Figure 9a shows the CE as a function of the polar angle in the PMT’s local coordinates for the NNVT MCP-PMT (red), NNVT HQE MCP-PMT (blue), and HPK dynode-PMT (violet). The CE curves are obtained by averaging the PDEs measured by the 7 LEDs along the zenith angles in the scanning station. The maximum CE values are set to 100% for the NNVT MCP-PMT and 93% for the HPK dynode-PMT, indicated by the electrostatic simulation results from NNVT and HPK private communications, respectively.
@@
where $i$ represents the index of the LED in the scanning station, and $w_{i}$ denotes the surface area weight of the $i$-th LED, as summarized in Table 2. This calculation is performed for each LPMT in the CD to ensure that the product of QE and CE is consistent with the measured PDE, as discussed in [21]. The QE spectral responses are also implemented in the simulation based on laboratory measurements, as shown in Fig. 9b. The NNVT MCP-PMT (red) and NNVT HQE MCP-PMT (blue) have identical QE spectral responses, while the HPK dynode-PMTs (violet) have a different response at higher wavelength ($\lambda>500~{}\mathrm{nm}$). It is assumed that LPMTs of the same type share the same CE curve and QE spectral response in the simulation.
Figure 9: (a) Angular dependence of collection efficiency for NNVT MCP-PMT (red), NNVT HQE MCP-PMT (blue), and HPK dynode-PMT (violet). (b) Wavelength dependence of quantum efficiency for NNVT MCP-PMT (red), NNVT HQE MCP-PMT (blue), and HPK dynode-PMT (violet).
The PMT optical model [22] takes into account the angle of incidence (AOI) dependence of the PDE, as well as the reflections on the photocathode and optical processes inside the PMTs. In this model, the PMT window is treated as a multi-layer optical stack, from outside to inside, consisting of water, PMT glass, anti-reflective coating, photocathode, and vacuum layers. The anti-reflective coating and photocathode are considered coherent layers due to their comparable thicknesses with the light wavelength. The model incorporates the light interference effect and the multiple reflections between adjacent boundaries using the transfer matrix method.
The refractive index $n$, extinction coefficient $k$, and thickness $d$ of the anti-reflective coating and photocathode are determined in the wavelength range of 390 nm to 500 nm by analyzing the reflectance data of NNVT MCP-PMT, NNVT HQE MCP-PMT, and HPK PMT immersed in the LAB liquid. The optical properties of the other components inside the PMTs, such as dynode (MCP), supporting structure, and aluminum film, are constrained by the QE data. These parameters are presented in [22] and are directly used in this work. With these inputs, the PMT optical model calculates both the reflectance and absorbance for a given photon and AOI, assuming uniformity for the anti-reflective coating and photocathode. The reflected photons are then propagated by GEANT4. The absorbance at a specific AOI is converted to the QE using the QE calculated above and absorbance information at normal incidence. This optical model was integrated into the detector simulation and used in this work. It has also been applied in the Daya Bay simulation, using the optical parameters obtained from the reflectance and QE data of an 8-inch Daya Bay PMT.
@@ 4.6.2 SPMTs
4.7 PE yield
In the updated detector simulation, the predicted PE yield of the JUNO detector is 1665 PE/MeV, calibrated using neutron capture events on hydrogen at the detector center. The PE yield is larger compared to the previously reported value of 1345 PE/MeV [30]. This enhancement can be attributed to three factors:
  • Improved PMT PDE: In previous studies, the LPMT’s PDE was assumed to be 27%. However, the LPMT mass testing has shown that the actual average PDE in CD is 30.1% [21]. This higher efficiency accounts for an $\sim$11increase in PE yield.
  • More realistic PMT optical model: Previous simulations used a simplified PMT optical model that neglected the angular dependence of PDE in water and PMT photocathode reflection. By incorporating a more realistic optical model [22], we observed an additional $\sim$8increase in PE yield.
  • Detector geometry updates: The detector geometry is updated based on the final mechanical design, with reflections on several detector components taken into account, leading to an approximate $\sim$3increase in PE yield.
Furthermore, the Cherenkov to total PE ratio is found to be 1.1% for 1 MeV positrons, and the detector simulation indicates a PE yield of 1785 PE/MeV for uniformly distributed neutron capture events on hydrogen.
@@ 5.1.1 LPMT contribution
When converting an SPE from the detector simulation into a pulse in the electronics simulation, the amplitude of the pulse is modeled by a combined distribution of a Gaussian and exponential function. The SPE charge resolution of each PMT, obtained from the PMT mass testing system [21], is directly used as the sigma of the Gaussian component in the simulation. The exponential component exhibits distinct behaviors between HPK LPMTs and NNVT LPMTs. For the former, the contribution from the exponential component is found to be 1% with an average amplitude of 1.1 PE, resulting in an SPE relative variance of 0.4. For the latter, the exponential component describes the large signal observed in the measured charge spectrum of MCP-PMT (blue line in Fig. 10a) with an average amplitude of 2.2 PE, resulting in an SPE relative variance of 0.7. The ratio of large signals in the SPE spectrum of MCP-PMT, as a function of the zenith angle on the photocathode, has been characterized by laboratory measurements, as shown in Fig. 10b. Good agreement between Monte Carlo simulation and experimental data is achieved, as illustrated in Fig. 10a.
@@
Figure 10: (a) The charge spectrum of NNVT MCP-PMT, measured with a pulsed light source, exhibits a long tail, which is described by an exponential distribution in the Monte Carlo simulation. Good agreement between data and Monte Carlo is achieved. (b) The ratio of the large signals depends on the positions on the photocathode, and is extracted from experimental data.
The transit time and TTS are also implemented for both HPK and NNVT LPMTs. As these quantities are only measured for approximately $25\%$ of LPMTs [21], a random sampling based on the distributions is used to obtain transit time and TTS values for other LPMTs in the CD. The transit time and TTS values are position-dependent on the photocathode, and this effect is also considered in the electronics simulation based on dedicated measurements made during the PMT mass testing.
The after-pulses of LPMTs are modeled based on the results published in [65]. As the earliest after-pulse component occurs at approximately 1 $\mu$s after the primary signals, the overlap with the signals is expected to be small and thus has a negligible impact on the energy resolution. The non-linearity effect of LPMTs is also taken into account in the electronics simulation. However, it primarily affects high-energy events rather than IBD ones.
@@ 5.1.2 Electronics contribution
In the LPMT electronics simulation, the SPE average waveform is obtained using the PMT testing data for both NNVT and HPK LPMTs. This waveform serves as a template in the electronics simulation to model the SPE response. The amplitude of the waveform is determined by the gain and SPE charge resolution discussed earlier.
To model the overshoot effect, the same model as that used in the Daya Bay experiment [66] is adopted, but with parameters tuned specifically for the JUNO experiment. The overshoot model is parameterized using an exponential function plus a Gaussian function. To capture the onset of the overshoot, the exponential function is multiplied by a Fermi function. The maximum amplitude of the overshoot is set to be 1% of the primary pulse [67].
White noise is applied to the raw waveform by sampling a Gaussian function with a mean of 10% of the SPE amplitude. This value is based on the requirements of the JUNO experiment. The digitization process is also modeled in the simulation, taking into account the effects of the front-end analog-to-digital converter (FADC) resolution, high-gain and low-gain amplifications of the SPE signal, and baseline offset. The simulation also considers the non-linearity effects of the electronics. However, it is anticipated that these effects will have a negligible impact on the energy resolution for positrons coming from IBD events.
Overall, the LPMT electronics simulation in the CD takes into account various aspects of the LPMT response, including the SPE waveform, overshoot, white noise, digitization, and non-linearity effects. These simulations aim to accurately model the electrical responses of the LPMTs and readout electronics and their impact on the energy resolution of the detector.
@@ 5.2 SPMT electronics simulation
The SPMT electronics simulation aims to accurately model the response of SPMTs and the readout electronics. For each SPMT, a DCR value is randomly sampled and assigned based on the DCR distribution obtained from the SPMT mass production measurements [26]. The mean values of the TTS (1.6 ns) and SPE charge resolution (33%) are shared by all SPMTs.
The key features of the CatiROC ASIC are implemented in the simulation based on the measurements in Ref. [34], including the two types of dead times: 70 ns trigger dead time and 9.3 $\mu$s digitization dead time. In cases where a hit is not triggered due to dead time, a fraction of its charge can still be calculated. This fraction is called the charge acceptance and is parameterized based on the time difference with the previous hit. In addition, the charge in picocoulombs is converted into an ADC unit, with two conversion coefficients that responsible for high gain (less than 10 PE) and low gain (more than 10 PE).
@@ 6.1 Reconstruction and calibration of PMT waveforms
As described in Sec. 5, the PMT together with the electronics system will convert the photons to waveforms for each LPMT and to the pairs of charge and time for each SPMT. For each LPMT waveform, the charge and hit time of every pulse are reconstructed using the deconvolution algorithm [68], which exhibits smaller charge non-linearity compared to other algorithms. Meanwhile, the hit time of the first pulse, referred to as the first hit time, is of particular importance. A dedicated algorithm with a linear fit of the rising edge of the first pulse is developed to achieve a more accurate first hit time and reduce its charge dependence. Given that the charge and hit time information of PMTs are the inputs to the vertex and energy reconstruction in Sec. 6.2, the performance of the waveform reconstruction will also contribute to the energy resolution. Moreover, the CD has approximately O(104) PMTs, and their characteristics may differ. Thus, the reconstructed charge and hit time of each PMT must be calibrated to account for the different PMT parameters, such as gain, TTS, relative PDE, and DCR. A comprehensive calibration strategy was developed in Ref. [30] to extract these parameters for all the PMTs and continuously monitor their time dependence.